### v2

# Were Residual Penalty and Neural Operators All We Needed for Solving Optimal Control Problems?

Abstract
Neural networks have been used to solve optimal control problems, typically by training neural networks using a combined loss function that considers data, differential equation residuals, and objective costs. We show that including cost functions in the training process is unnecessary, advocating for a simpler architecture and streamlined approach by decoupling the optimal control problem from the training process. Thus, our work shows that a simple neural operator architecture, such as DeepONet, coupled with an unconstrained optimization routine, can solve multiple optimal control problems with a single physics-informed training phase and a subsequent optimization phase. We achieve this by adding a penalty term based on the differential equation residual to the cost function and computing gradients with respect to the control using automatic differentiation through the trained neural operator within an iterative optimization routine. Our results show acceptable accuracy for practical applications and potential computational savings for more complex and higher-dimensional problems.
This work has been submitted to IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I INTRODUCTION
An optimal control problem is an optimization problem in which the system dynamics are described by differential equations, either ordinary differential equations (ODEs) or partial differential equations (PDEs), that explicitly depend on a control input. These problems arise across a wide range of practical applications, including power plant operation [1], flight trajectory optimization [2], control of biological systems [3], and economic decision-making in areas such as investment and trading [4].
Traditional numerical approaches to optimal control problems are computationally intensive, making neural networks an attractive alternative. As such, recent studies have explored physics-informed neural networks (PINNs) and neural operators (NOs) for solving optimal control problems [5, 6, 7, 8, 9], particularly when the constraints are complex, nonlinear PDEs. Existing neural network models are typically trained for a single task, requiring retraining if problem parameters or cost functions change. For example, in a tracking problem, the tracked target may change, leading to a corresponding change in the cost function. In flight control, the objective may vary depending on the mission phase, such as achieving the fastest climb, the most economical cruise, or the safest landing, each defining a different optimal control problem. Similarly, operating a power plant involves distinct objectives during startup, steady-state operation, and shutdown, resulting in various formulations of the cost function depending on the operational context.
To address this limitation, we propose a method that fully decouples neural network training from the optimal control problem, eliminating the need for retraining when cost functions or parameters change. Our approach trains a single NO in advance, allowing it to solve multiple optimal control problems efficiently. Additionally, we argue that solving a complex optimal control problem is fundamentally equivalent to training a well-generalizable physics-informed NO, making specialized architectures unnecessary. In line with Occam’s razor, we advocate for simpler and less complex architectures, as our results demonstrate that such models are sufficient for solving a wide range of control problems without the need for elaborate or highly specialized designs. In addition, our approach broadens the original scope of NOs beyond their typical use as solvers for differential equations, thus showcasing a valuable discovery of their capabilities.
Our method employs a direct optimization approach, integrating automatic differentiation with nonlinear programming to efficiently compute control gradients while enforcing differential equation constraints through residual penalties. A schematic representation of our approach is shown in Figure 1. We focus on the DeepONet architecture, although our approach is not limited to this architecture, leaving exploration of alternatives for future work.
Our main contributions are as follows. We demonstrate that a single NO can be reused with our method across multiple optimal control problems, eliminating the need for retraining, thereby enabling their use in practical settings. We further demonstrate that training a NO alone is sufficient for solving optimal control problems, reducing the need for complex architectural modifications. This suggests that, in many cases, specialized control networks may not be necessary, offering a more streamlined approach for using neural networks with optimal control problems. Our approach is both simple and powerful, showcasing the versatility of NOs in solving a broad range of problems beyond their original task. To validate its applicability and efficiency, we benchmark our method against established reference solutions for both ODE and PDE control problems.
The rest of this paper is structured as follows: Section II reviews related work of neural networks in optimal control. Section III introduces our direct method. Section IV presents experimental results on two ODE- and two PDE-constrained optimal control problems with two different cost functions for each ODE or PDE. Section VI concludes the study.
Figure 1: Schematic of the proposed framework. A neural operator (DeepONet) maps the control input $u$ and space–time coordinates $(x,t)$ into the system state $y$. The state is used to evaluate a penalized objective $J_{\mu}(u)$, which combines the original cost functional with physics residuals and regularization terms. Gradients $dJ_{\mu}(u)/du$ are computed via automatic differentiation, and the control is updated through an optimization step. This process iterates until convergence, yielding the optimal control.
II RELATED WORK
In recent years, PINNs have been widely used to solve differential equations [10, 11, 12, 13, 14]. An extension of PINNs for solving optimization problems was performed by Lu et al. [15], where the authors used a PINN and a control network for topology optimization for inverse design problems. They trained a neural network on a common loss function consisting of the residual of the differential equation and the cost function of the topology optimization problem. Molawi et al. [5] extended the idea of a control neural network and a PINN to optimal control problems. Song et al. [9] split the control term into two variables, dividing the training loss into a differential equation term and a regularizing term describing the cost, and used an alternating direction method of multipliers for training the neural networks. Similarly, [16, 7, 8] have explored the use of two (or more) neural network models, PINNs and a control neural network, to solve optimal control problems.
Another common approach is to utilize the adjoint state equation, which arises from Pontryagin’s Minimum Principle [17] together with PINNs. Employing Pontryagin’s Minimum Principle transforms the optimal control problem into a set of differential equations defined by the Hamiltonian of the system and adjoint state equations. Yin et al.[18] used a direct adjoint looping approach to solve the optimal control problem by using two separate neural networks to solve it, one for the differential equation and one for the adjoint equation that arises from Pontryagin’s Minimum Principle. Schiassi et al. [19] also used Pontryagin’s Minimum Principle and trained a PINN for solving the two-point boundary value problem for a quadratic cost function directly. Demo et al. [16] used a PINN and a control network in sequence, where the output of the PINN was given as input to the control network, which solved for the adjoint equation and derived the optimal control. These methods showed promising results on their respective test problems, but are only capable of solving the particular optimal control problem (i.e., with fixed cost function parameterization) for which the neural networks were trained.
Physics-informed NOs have also been applied to solve optimal control problems [20, 21, 22]. Hwang et al. [23] showed how a modified DeepONet can be used to solve optimal control problems by first training a model and thereafter searching for the optimal control by using nonlinear programming and unconstrained optimization routines, where the gradients of the cost function were calculated by automatic differentiation with respect to the control. The approach in [23] is similar to our proposed idea, however, the authors used an autoencoder to reconstruct the control function before re-feeding it to the NO model in the optimization routine’s iterations. In contrast, we show how this can be done without an autoencoder by properly penalizing the cost function with the physics residual, thus avoiding any increased complexity in the network architecture and associated more computationally expensive training process.
III METHODOLOGY
III-A Neural Operators
Consider the following general differential equation, which may represent either an ODE or a PDE:
$\mathcal{D}(\mathbf{y})(\mathbf{x},t)$ $=\bm{f}(\mathbf{y}(\mathbf{x},t),\mathbf{u}(\mathbf{x},t),\mathbf{x},t),$ (1)
$\mathcal{C}(\mathbf{y})$ $=\bm{g}(\mathbf{x},t),$
where $\mathbf{x}\in\Omega\subset\mathbb{R}^{d}$ denotes the spatial coordinate (with $d=0$ in the ODE case), and $t\in[0,T]$ is time. Here, $\mathbf{y}(\mathbf{x},t)\in\mathbb{R}^{n}$ is the state, $\mathbf{u}\in\mathcal{U}\subset C^{\infty}(\Omega\times[0,T];\mathbb{R}^{m})$ is the control function, $\mathcal{C}(\mathbf{y})=\bm{c}$ enforces the initial condition on $\Omega\times\{0\}$ and the boundary condition on $\partial\Omega\times(0,T]$, $\mathcal{D}$ is a differential operator representing the system dynamics, and $\bm{f}$ is a nonlinear function describing inputs and sources. To ease the notation, we omit the arguments, for example, by denoting $\mathbf{u}(\mathbf{x},t):=\mathbf{u}$, only including them when the context requires. In addition, boldfaced variables represent vectors, while nonboldfaced variables are scalars.
We define $\mathbf{y}$ as the solution of (1). We seek to find an operator $\mathcal{G}:\mathcal{U}\times\Omega\times[0,T]\rightarrow\mathbb{R}^{n}$ that solves the differential equation (1) for any given function $\mathbf{u}$. Hence, we write the operator as
$\mathcal{G}(\mathbf{u})(\mathbf{x},t)=\mathbf{y}.$ (2)
A Neural Operator approximates operator $\mathcal{G}$ with a neural network. Lu et al. [24] presented an architecture, the Deep Operator Network (DeepONet), based on the universal approximation theorem for operators [25]. The basic architecture consists of two parallel networks called branch and trunk. The branch network receives as input the function $\mathbf{u}$ discretized into $m$ points called sensor locations, thus denoted by $\{\mathbf{u}_{k}\}_{k=0}^{m-1}\in\mathbb{R}^{n\times m}$, where curly braces indicate a vector or sequence of discretized values. Respectively, the trunk network receives discrete space–time coordinates $(\mathbf{x}_{i},t_{i})$, where the single index $i$ enumerates the discretization points in both space and time. Both the trunk and branch networks are fully connected networks in their simplest form. At the end of the DeepONet, the outputs of the trunk and branch networks are combined with a dot product. Generally, a NO seeks to simultaneously approximate the operator $\mathcal{G}$ and solve the differential equation (1). We define the NO $\mathcal{G}_{\theta}:\mathbb{R}^{n\times m}\times\Omega\times[0,T]\rightarrow\mathbb{R}^{n}$ with parameters $\theta$ and at a discrete space–time point $(\mathbf{x}_{i},t_{i})$ as
$\mathcal{G}_{\theta}(\{\mathbf{u}_{k}\}_{k=0}^{m-1})(\mathbf{x}_{i},t_{i})=\tilde{\mathbf{y}}_{i},$ (3)
where $\tilde{\mathbf{y}}\approx\mathbf{y}$. In practice, multiple coordinate points $\{(\mathbf{x}_{i},t_{i})\}$ can be input as a batch for efficient evaluation, but we omit this in the notation for clarity. For more details, we refer the reader to [24, 26, 27].
Training the NO can be performed in a purely data-driven fashion using pairs of input functions and corresponding solutions (e.g., minimizing $\|\tilde{\mathbf{y}}-\mathbf{y}\|$), or in a physics-informed way. In the latter, the loss function incorporates the residual of the governing differential equation (1) and boundary/initial conditions. This enforces physical consistency by penalizing violations of $\mathcal{D}(\tilde{\mathbf{y}})-\bm{f}=0$ and $\mathcal{C}(\tilde{\mathbf{y}})-\bm{g}=0$, enabling training even with no data while embedding the dynamics directly into the learning process. We define the discrete differential residual function $\mathcal{R}_{i}=\mathcal{R}_{i}(\bm{f},\tilde{\mathbf{y}}_{i},\mathbf{u}_{i},\mathbf{x}_{i},t_{i})$ as
$\mathcal{R}_{i}=\left\|\mathcal{D}(\tilde{\mathbf{y}}_{i})-\bm{f}(\tilde{\mathbf{y}}_{i},\mathbf{u}_{i},\mathbf{x}_{i},t_{i})\right\|_{2}^{2},$ (4)
and the constraint residual, which includes the boundary and initial conditions, as
$\mathcal{R}^{c}_{i}=\left\|\mathcal{C}(\tilde{\mathbf{y}}_{i})-\bm{g}(\mathbf{x}_{i},t_{i})\right\|_{2}^{2}$ (5)
A central element in this way of training the NO is the differential operation $\mathcal{D}(\tilde{\mathbf{y}}_{i})$, which can be easily computed with automatic differentiation at the discrete points. Thus, we summarize the process of training a physics-informed NO by solving an optimization problem
$\min_{\theta}\frac{1}{|S|}\sum_{i\in S}\mathcal{R}_{i}+\frac{1}{|\partial S|}\sum_{j\in\partial S}\mathcal{R}^{c}_{j},$ (6)
where $S$ denotes the set of inner collocation points and $\partial S$ the boundary and initial collocation points. For more technical details, implementations and variants of physics-informed training we refer to [10, 11, 26].
III-B A Direct Method for Optimal Control Problems
A typical optimal control problem $\min_{\mathbf{u}}J(\mathbf{u})$ can be defined as
$\min_{\mathbf{u}}\quad\int_{0}^{T}\int_{\Omega}$ ${L(\mathbf{y},\mathbf{u},t)\mathbf{dx}dt}+E(\mathbf{y}(\mathbf{x},T),\mathbf{u}(\mathbf{x},T))$ (7)
s.t. $\quad\mathcal{D}(\mathbf{y})=\bm{f}(\mathbf{y},\mathbf{u},\mathbf{x},t)$
$\quad\mathcal{C}(\mathbf{y})=\bm{g}(\mathbf{x},t),$
where $L$ is the running cost and is a smooth function with the mapping $L:\mathbb{R}^{n}\times\mathcal{U}\times\Omega\times[0,T]\rightarrow\mathbb{R}$, and $E$ is the terminal cost with mapping $E:\mathbb{R}^{n}\times\mathcal{U}\times\Omega\rightarrow\mathbb{R}$. The constraint equation $\mathcal{D}(\mathbf{y})=\bm{f}(\mathbf{y},\mathbf{u},\mathbf{x},t)$ is a set of differential equations with the boundary and initial conditions $\mathcal{C}(\mathbf{y})=g(\mathbf{x},t)$.
For solving (7), one can resort to the indirect methods using Pontryagin’s Minimum Principle [17] for ODEs and the adjoint method [28] for PDEs and transfer the problem to another set of differential equations which need to be solved. Another approach is the direct method, which discretizes the optimal control problem and treats it as a nonlinear programming (NLP) problem [29]. This can be achieved by discretizing the space–time domain into grid points indexed by $i=0,\dots,I$. For convenience, we denote the index set as $\mathcal{I}:=\{0,1,\dots,I\}$. Here, $\mathcal{I}$ denotes the union of all grid indices, i.e., interior points $\mathcal{I}_{\mathrm{int}}$, boundary points $\mathcal{I}_{\mathrm{b}}$, initial points $\mathcal{I}_{0}$, and terminal points $\mathcal{I}_{T}$. The discrete approximations of the functions are denoted as $\mathbf{y}_{i}\approx\mathbf{y}(\mathbf{x},t)$ and $\mathbf{u}_{i}\approx\mathbf{u}(\mathbf{x},t)$ at the $i$-th grid point. This yields the discrete optimization problem $\min\bar{J}(\{\mathbf{u}_{i}\}_{i\in\mathcal{I}})$ defined as
$\min_{\{\mathbf{u}_{i}\}_{i\in\mathcal{I}}}$ $\sum_{i\in\mathcal{I}}L(\mathbf{y}_{i},\mathbf{u}_{i},\mathbf{x}_{i},t_{i})\,\Delta w_{i}+E(\mathbf{y}_{T},\mathbf{u}_{T}),$ (8)
s.t. $\mathcal{D}(\mathbf{y}_{i})=\bm{f}(\mathbf{y}_{i},\mathbf{u}_{i},\mathbf{x}_{i},t_{i}),\quad\forall i\in\mathcal{I},$
$\mathcal{C}(\mathbf{y}_{i})=\bm{g}(\mathbf{x}_{i},t_{i}),\quad\forall i\in\mathcal{I}_{\mathrm{b}}\cup\mathcal{I}_{0},$
where $\Delta w_{i}$ are quadrature weights from the chosen integration rule such as Gaussian quadrature or trapezoidal integration. The notation $\mathbf{y}_{T},\mathbf{u}_{T}$ refers to the collections of discrete values at the terminal time, i.e. $\{\mathbf{y}_{i},\mathbf{u}_{i}:i\in\mathcal{I}_{T}\}$. The single index $i$ implicitly enumerates both space and time discretization points. Equation (8) is now an NLP problem.
In general, solving NLP problems require the computation of gradients of the cost function $\bar{J}(\mathbf{u}_{i})$ with respect to $\mathbf{u}_{i}$, i.e., $d\bar{J}(\mathbf{u}_{i})/d\mathbf{u}_{i}$. Thus, solving a discretized optimal control problem directly is not trivial as it requires the computation of the partial derivative of the state $\frac{\partial\mathbf{y}_{i}}{\partial\mathbf{u}_{i}}$ due to the chain rule. Fortunately, neural networks are a suitable surrogate in settings where one needs to compute gradients of the cost and the state, as neural networks are differentiable by automatic differentiation with respect to the input. Next, we show how this can be used to our benefit.
III-C Neural Operators for Direct Methods
The discretized optimal control problem (8) can be solved as an NLP problem with NOs acting as a surrogate for the differential equation constraints. That is, we replace the differential equations in the constraints of (8) with our pre-trained NO. To ensure that the integrals in the cost function are evaluated without the need for interpolation, the coordinate discretization points $(\mathbf{x}_{i},t_{i})$ are chosen to coincide with the sensor locations $\mathbf{u}_{i}=\mathbf{u}(\mathbf{x}_{i},t_{i})$. This leads to the following formulation:
$\min_{\{\mathbf{u}_{i}\}_{i\in\mathcal{I}}}$ $\sum_{i\in\mathcal{I}}L(\tilde{\mathbf{y}}_{i},\mathbf{u}_{i},\mathbf{x}_{i},t_{i})\,\Delta w_{i}+E(\tilde{\mathbf{y}}_{T},\mathbf{u}_{T}),$ (9)
s.t. $\quad\mathcal{G}_{\theta}(\{\mathbf{u}_{i}\}_{i\in\mathcal{I}})(\mathbf{x}_{i},t_{i})=\tilde{\mathbf{y}}_{i},\quad\forall i\in\mathcal{I}.$
Since $\mathcal{G_{\theta}}$ is a neural network (a NO), the output of the neural network can be differentiated with respect to the input, thus, an optimization routine for NLP problems, such as gradient descent (GD), can be employed to solve the problem. For that, we require only two things. First, the NO must be trained, generalizable, and expressive enough to have a wide enough search space. This can be ensured in practice by making the number of functions in the NO training set large enough. Second, since we want to optimize on the control (input) $\mathbf{u}$ and repeatedly update $\mathbf{u}$, the NO can produce solutions $\mathbf{y}$ that do not approximate the solution to the differential equation. To address this, the cost function must be properly penalized with the residual of the differential equation, ensuring that the control remains in the known search space and satisfies the differential equation. Without properly penalizing with the residual, the computed gradients are noisy and the method fails to find the solution as shown in Figure 2. Thus, we penalize the cost function (9) with the differential equation residual (4), obtaining
$\min_{\{\mathbf{u}_{i}\}_{i\in\mathcal{I}}}$ $\bar{J}_{\mu}(\{\mathbf{u}_{i}\}_{i\in\mathcal{I}})=\min_{\{\mathbf{u}_{i}\}_{i\in\mathcal{I}}}\sum_{i\in\mathcal{I}}L(\mathbf{y}_{i},\mathbf{u}_{i},\mathbf{x}_{i},t_{i})\,\Delta w_{i}$ (10)
$+$ $E(\mathbf{y}_{T},\mathbf{u}_{T})+\frac{\mu}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathcal{R}_{i}+\lambda\operatorname{Reg}(\{\mathbf{u}_{i}\}_{i\in\mathcal{I}})$
$\text{s.t.}\quad\mathcal{G}_{\theta}(\{\mathbf{u}_{i}\}_{i\in\mathcal{I}})(\mathbf{x}_{i},t_{i})=\mathbf{y}_{i}\quad\forall i\in\mathcal{I},$
where $\mu\in\mathbb{R}_{+}$ is a penalty factor and we have added an optional regularization operator $\operatorname{Reg}$ with the regularization parameter $\lambda\in\mathbb{R}_{+}$. The regularization is optional but can provide smoother solutions and reduce noise, for example with a Tikhonov regularization $\|\mathbf{u}\|_{2}^{2}$ or the $H_{2}$-seminorm regularization $\|\nabla^{2}u\|_{2}^{2}$. Given a sufficiently large penalty factor $\mu$, it enforces the optimization routine to remain within the domain where the residuals of the differential equation remain small. The penalized cost function $\bar{J}_{\mu}$ is differentiable with respect to the control $\mathbf{u}$ since the differential equation was replaced with the NO $\mathcal{G}_{\theta}$. The differentiability of the cost function with respect to $\mathbf{u}$ allows us to use any optimization routine that utilizes gradients of the penalized cost function. We summarize the solution process in Algorithm 1, where $\mathbf{u}_{i}\leftarrow\Phi(\mathbf{u}_{i},\frac{d\tilde{J}_{\mu}}{d\mathbf{u}_{i}})$ represents the current solution update step performed (e.g., a step in the gradient direction for GD) within the optimization routine.
Figure 2: Control solutions for Poisson’s problem (15) and cost (16a). (Left) NO direct method with residual penalty $\mathcal{R}$ and (right) without penalty.
Algorithm 1 Gradient-based Optimization using a Neural Operator
1:Input: Pre-trained Neural Operator $\mathcal{G}_{\theta}$; discretized cost $\bar{J}(\mathbf{u})$
2:Output: Optimal control $\mathbf{u}^{*}$
3:Initialize: $\{\mathbf{u}_{i}\}_{i\in\mathcal{I}}$ (e.g., zeros); define $\mathbf{x}_{i}$ and $t_{i}$ to match with $\mathbf{u}_{i}\equiv\mathbf{u}(\mathbf{x}_{i},t_{i})$
4:while not converged do
5:  $\tilde{\mathbf{y}}_{i}=\mathcal{G}_{\theta}(\{\mathbf{u}_{i}\}_{i\in\mathcal{I}})(\mathbf{x}_{i},t_{i})\quad\forall i\in\mathcal{I}$
6:  $\mathcal{R}_{i}=\left\|\mathcal{D}(\tilde{\mathbf{y}}_{i})-\bm{f}(\tilde{\mathbf{y}}_{i},\mathbf{u}_{i},\mathbf{x}_{i},t_{i})\right\|_{2}^{2}\quad\forall i\in\mathcal{I}$
7:  $\bar{J}_{\mu}=\bar{J}(\{\mathbf{u}_{i}\}_{i\in\mathcal{I}})+\frac{\mu}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathcal{R}_{i}+\lambda\operatorname{Reg}(\{\mathbf{u}_{i}\}_{i\in\mathcal{I}})$
8:  Compute $\frac{d\bar{J}_{\mu}}{d\mathbf{u}_{i}}\quad\forall i\in\mathcal{I}$
9:  $\mathbf{u}_{i}\leftarrow\Phi\!\left(\mathbf{u}_{i},\frac{d\bar{J}_{\mu}}{d\mathbf{u}_{i}}\right)\ \forall i\in\mathcal{I}$
10:end while
11:return $\mathbf{u}^{*}\leftarrow\{\mathbf{u}_{i}\}_{i\in\mathcal{I}}$
IV EXPERIMENTAL SETUP
We conduct experiments by training DeepONet models for optimal control problems with varying cost functions, showcasing our proposed method’s re-usability, how it extends the original scope of NOs, and its architectural simplicity. The models are used for solving the discretized optimal control problem (10) by our method in Algorithm 1. We also solve the problems with a reference method, to give a comparison in both computing times and solution accuracy.
IV-A Experimental Problems
We study the following dynamical systems. For each model, we list two different cost functions, and solve the arising optimal control problems with our method, demonstrating how the trained neural operator can be reused in our setting.
  1. 1.
    Nonlinear ODE:
    $\frac{dy}{dt}(t)$ $=\tfrac{5}{2}\Big(-y(t)+y(t)u(t)-u(t)^{2}\Big),$ (11)
    $y(0)$ $=1,\quad t\in[0,1].$
    This dynamical system has been used as a benchmark by [30] and [31] for solving optimal control problems. We choose to start with this easy problem to demonstrate that the method works for both terminal cost functions and continuous cost functions, hence
    $J_{\mathrm{ODE},1}(u)$ $=-y(1)$ (12a)
    $J_{\mathrm{ODE},2}(u)$ $=\int_{0}^{1}(y(t)^{2}+u(t)^{2})dt$ (12b)
  2. 2.
    SIR-model:
    $\frac{dS(t)}{dt}$ $=-\beta(1-u(t))S(t)I(t)-v(t)S(t),$ (13)
    $\frac{dI(t)}{dt}$ $=\beta(1-u(t))S(t)I(t)-\gamma I(t),$
    $\frac{dR(t)}{dt}$ $=\gamma I(t)+v(t)S(t),$
    where $0,\displaystyle S(0)=98,I(0)=02,,R(0)=0$,
[…]
    This epidemic control model follows the standard Susceptible-Infected-Recovered (SIR) formulation with an additional vaccination control $v(t)$, similar to the setting in [32], and a control of contacts, $u(t)$. The compartmental models are widely studied as an application of optimal control in epidemiology. We choose this problem to demonstrate that the method works for multidimensional ODEs and for multiple control functions. We choose two cost functions, the first one considers vaccination and control planning for minimizing costs. The second incorporates healthcare capacity constraints by adding a penalty on the infected population whenever it exceeds 20%,
    $J_{\mathrm{SIR},1}(u,v)$ $=\int_{0}^{10}\!\Big(I^{2}+0.2u^{2}+5v^{2}\Big)\,dt,$ (14a)
    $J_{\mathrm{SIR},2}(u,v)$ $=\int_{0}^{10}\!\Big(0.1u^{2}+v^{2}+10\bigl[I-0.2\bigr]_{+}^{2}\Big)\,dt,$ (14b)
  3. 3.
    Poisson’s problem:
    $\nabla^{2}y(\mathbf{x})$ $=u(\mathbf{x}),\qquad\mathbf{x}\in\Omega,$ (15)
    $y(\mathbf{x})$ $=0,\qquad\mathbf{x}\in\partial\Omega,$
    $\Omega$ $=[0,1]\times[0,1].$
    This problem is a test case from [23] and reflects targeting a heat profile with a given source. We choose this problem as it demonstrates that our method works for time-independent PDEs and with multidimensional controls. We choose both cost functions to be of tracking type, where the first cost function follows a state that was generated with $u_{ref,1}(x_{1},x_{2})=\sin(\pi x_{1})\sin(\pi x_{2})$, which is a scaled version from [23]. The second tracking object is a solution obtained by a control generated by a Gaussian random field (GRF) with a lengthscale of $l=1.0$.
    $J_{\mathrm{Pois},1}(u)$ $=\int_{\Omega}\!\big(y-y_{ref,1}\big)^{2}\,d\mathbf{x},$ (16a)
    $J_{\mathrm{Pois},2}(u)$ $=\int_{\Omega}\!\big(y-y_{ref,2}\big)^{2}\,d\mathbf{x}$ (16b)
  4. 4.
    Forced Burger’s equation:
[…]
    $y(x,0)$ $=0,\quad y|_{\partial\Omega}=0,$
    $(x,t)\in\Omega$ $\times[0,T],\ \Omega=[0,1].$
    We choose this problem in order to demonstrate that our method works for nonlinear time-dependent PDEs as well as multidimensional controls. A variant of this problem was given in [23]. We choose two cost functions of tracking type, where the tracked references are solutions of controls generated by a GRF with lengthscale of $l=1.0$ for $J_{\mathrm{Burg},1}$ and $l=0.5$ for $J_{\mathrm{Burg},2}$.
    $J_{\mathrm{Burg},1}(u)$ $=\int_{0}^{T}\!\!\int_{\Omega}\big(y-y_{ref,1}\big)^{2}\,dx\,dt,$ (18a)
    $J_{\mathrm{Burg},2}(u)$ $=\int_{0}^{T}\!\!\int_{\Omega}\big(y-y_{ref,2}\big)^{2}\,dx\,dt.$ (18b)
IV-B Network Architectures and Training.
We employed the DeepONet architecture, using a modified version with residual connections in the fully connected networks as described by Wang et al. [27]. All models used the $\tanh$ activation function. For training, we generated input functions from two families: GRFs with lengthscales in $[0.05,1]$ and polynomials up to degree 5. The architecture and training settings are summarized in Table I, and all input functions were normalized to the ranges reported therein.
Training was physics-informed, with the loss consisting of the differential residual (4) and boundary residual (5). Residuals were computed at $1000$ randomly sampled space–time points using automatic differentiation. Optimization was performed in JAX with Adam ($10^{5}$ iterations, learning rate $10^{-4}$, batch-size $1000$) .
TABLE I: DeepONet model configurations.
Model $m$ Input Range Branch (W $\times$ D) Trunk (W $\times$ D)
Nonlinear ODE 100 [1.5,1.5][-1.5,1.5] $200\times 6$ $200\times 6$
SIR-model 200 [1,1][-1,1] $200\times 6$ $200\times 6$
Poisson $[32,32]$ [2,2][-2,2] $1024\times 3$ $200\times 5$
Forced Burgers’ eq. $[32,32]$ [2,2][-2,2] $1024\times 3$ $400\times 6$
IV-C Optimization and optimization parameters
We solve the different problems by formulating them as described in (10) and employing the solution method described in Algorithm 1. We summarize the parameters in Table II where the penalty factor $\mu$ and regularization factor $\lambda$ are defined in (10) as well as the regularization type. Finally, we use the parameter $\gamma$ as the step size for the updating step $\mathbf{u}_{i}\leftarrow\Phi(\mathbf{u}_{i},\frac{d\tilde{J}_{\mu}}{d\mathbf{u}_{i}})$.
TABLE II: Hyperparameters and regularization settings used in the experiments.
Problem $\gamma$ $\mu$ Reg. Type $\lambda$ $n_{iter}$
$J_{\mathrm{ODE},1}(u)$ 0.002 50 $H^{2}$ 0.1 2000
$J_{\mathrm{ODE},2}(u)$ 0.002 50 $H^{2}$ 1.0 2000
$J_{\mathrm{SIR},1}(u,v)$ 0.01 100 $H^{2}$ 0.02 2000
$J_{\mathrm{SIR},2}(u,v)$ 0.01 100 $H^{2}$ 0.02 2000
$J_{\mathrm{Pois},1}(u)$ 0.01 10 2000
$J_{\mathrm{Pois},2}(u)$ 0.01 10 2000
$J_{\mathrm{Burg},1}(u)$ 0.01 0.01 2000
$J_{\mathrm{Burg},2}(u)$ 0.01 0.01 2000
IV-D Reference Solutions
For comparing accuracy and computing times we solve all problems with a reference method. For the ODE problems, nonlinear ODE and SIR-model, we use an indirect approach with a shooting method to solve the optimal control problems, implemented in CasADi [33]. For the PDE problems, Poisson’s problem and the forced Burgers’ equation, we employ the adjoint method, using FEniCSx with DOLFINx [34] as the solver for both the differential equation and its adjoint. We use the same grid-size, collocation points (sensor locations), and update step size as in our neural operator approach. Our proposed neural-operator-based optimization is implemented in JAX [35] and executed on the GPU, whereas all reference solutions are computed on the CPU. We ran the problems on a system with 12th Gen Intel(R) Core(TM) i5-12600K, 32GB RAM and a NVIDIA RTX A2000 12GB.
V RESULTS
We present our results visually in the following subsections and summarize solution time, time per iteration, mean square error (MSE) and standard deviation (SD) of the control, and error in the final objective at the end of this section
V-A Nonlinear ODE
The state trajectories, control, and convergence plots for the nonlinear ODE problem are shown in Figure 3. While both our method and the reference solver converge to the same solution, our method exhibits oscillations and struggles to capture the initial control value accurately for the second cost function $J_{ODE,2}$. We attribute this to the explicit enforcement of initial conditions during neural operator training, which prevents accurate gradient computation at the initial time. The reference method shows a much faster convergence rate.
Figure 3: Comparison between our method (solid line) and the reference solution (dotted line). The top row corresponds to Problem 1 ($J_{\mathrm{ODE},1}$) and the bottom row to Problem 2 ($J_{\mathrm{ODE},2}$). Each row shows the state trajectory $y(t)$, the control $u(t)$, and the cost reduction across iterations.
Figure 4: Comparison between our method (solid line) and the reference solution (dotted line). The top row corresponds to Problem 1 ($J_{\mathrm{SIR},1}$) and the bottom row to Problem 2 ($J_{\mathrm{SIR},2}$). Each row shows the states $S(t),I(t),$ and $R(t)$, the control $u(t)$, and the cost reduction across iterations.
V-B SIR-model
The state trajectories, control, and convergence plots for the SIR model are shown in Figure 4. Both our method and the reference solver converge to the same overall solution, though some discrepancies remain between the two. Our method shows minor oscillations towards the end time. Here, also, the reference method shows a much faster convergence rate.
V-C Poisson’s Problem
The control, absolute control errors, and convergence plots for the Poisson problem are shown in Figure 5. For the smooth tracking function of $y_{ref,1}$, the reference method achieves significantly higher accuracy than our approach, while our method struggles in particular near the corners. For the second cost functional, $J_{\mathrm{Pois},2}$, even the reference solution shows difficulties at the domain boundaries. In terms of convergence, the reference method is faster for $J_{\mathrm{Pois},1}$, whereas both methods perform comparably for $J_{\mathrm{Pois},2}$.
Figure 5: Comparison between our method (left) and the reference solution (middle). The top block corresponds to Problem 1 ($J_{\mathrm{Pois},1}$) and the bottom block to Problem 2 ($J_{\mathrm{Pois},2}$).
Figure 6: Comparison between our method (left) and the reference solution (middle). The top block corresponds to Problem 1 ($J_{\mathrm{Burg},1}$) and the bottom block to Problem 2 ($J_{\mathrm{Burg},2}$).
TABLE III: Solution metrics. Best per column in bold (lower is better).
Problem Solution time [s] Time/iteration [s] Control MSE (± SD) Objective error
Ours Ref. Ours Ref. Ours Ref. Ours Ref.
$J_{\mathrm{ODE},1}(u)$ 2.51 0.010 1.3e-03 1.4e-03 3.2e-05 (± 5.7e-03) 1.2e-05 (± 9.2e-04) -2.0e-03 1.2e-06
$J_{\mathrm{ODE},2}(u)$ 2.53 8.9e-03 1.3e-03 1.3e-03 7.3e-04 (± 0.027) 1.5e-04 (± 0.011) 1.9e-05 8.7e-05
$J_{\mathrm{SIR},1}(u)$ 8.03 14.64 4.0e-03 0.98 1.5e-04 (± 0.011) 6.9e-06 (± 2.4e-03) 2.3e-03 5.4e-06
$J_{\mathrm{SIR},2}(u)$ 8.11 17.98 4.1e-03 1.00 4.8e-04 (± 0.022) 1.5e-05 (± 3.5e-03) 2.3e-03 2.7e-06
$J_{\mathrm{Pois},1}(u)$ 8.39 4.73 4.2e-03 2.4e-03 6.8e-03 (± 0.079) 9.0e-04 (± 0.028) 9.2e-05 3.2e-10
$J_{\mathrm{Pois},2}(u)$ 8.43 4.97 4.2e-03 2.5e-03 0.029 (± 0.17) 0.030 (± 0.17) 3.5e-05 1.1e-05
$J_{\mathrm{Burg},1}(u)$ 15.60 369.51 7.8e-03 0.18 0.043 (± 0.21) 0.11 (± 0.33) 2.4e-05 2.7e-05
$J_{\mathrm{Burg},2}(u)$ 15.95 357.62 8.0e-03 0.18 0.048 (± 0.22) 0.097 (± 0.31) 2.6e-05 2.1e-05
V-D Forced Burger’s Equation
The control, absolute control errors, and convergence plots for the forced Burger’s equation problem are shown in Figure 6. Both methods struggle to recover the true control near the domain edges, while our method additionally exhibits higher-frequency noise in the control error at interior points. The convergence rates of the two methods are similar.
V-E Solution Metrics
We summarize solution times and error metrics in Table III. The reference method performs better for the simple nonlinear ODE, both in time and error, while for the SIR model the neural operator achieves comparable accuracy and better total solution times. For the Poisson problem, the neural operator also shows competitive performance, but the reference (adjoint) method is faster, since solving Poisson’s equation reduces to a linear system that can be handled efficiently. In contrast, for the forced Burgers problem, our neural operator approach achieves substantially faster solution times, about two orders of magnitude per iteration, while maintaining similar accuracy. This advantage arises because the PDE is time-dependent, thus, each optimization step with the adjoint method requires marching forward in time across all spatial coordinates, which is less efficient than solving a linear system. We acknowledge that the adjoint method might improve with a more careful choice of step size for gradient updates, potentially reducing the number of iterations. Nevertheless, its per-iteration cost would remain significantly higher.
VI CONCLUSIONS
We demonstrated that a physics-informed DeepONet model can be directly applied to solve optimal control problems without requiring any information about the cost functions during training. This significantly enhances the practical applicability of neural operators and reduces the architectural complexity typically associated with solving optimal control problems using neural networks. Our results further show that neural operators can be used beyond their original role as differential equation solvers. Unlike conventional neural network approaches, we did not employ dedicated control networks or auxiliary components to explicitly construct the solution. Instead, the trained DeepONet was integrated into an optimization routine, where the optimal control problem was discretized and a differential-equation residual was added to the cost as a penalty, ensuring that the control remained within the solution space during iterative optimization. The method achieved competitive speed compared to the reference solver in the SIR model and is expected to outperform it for higher-dimensional ODE problems. For the time-dependent PDE (forced Burgers’ equation), our approach yielded iteration times up to 20× faster than the reference method, highlighting its potential for more complex time-dependent PDE-control problems. A clear advantage of our approach is that, for PDE-control, we never need to implement a PDE solver or solution scheme within the optimization loop itself; instead, the physics-informed training of the neural operator embeds the dynamics directly.
While our study demonstrated the potential of physics-informed DeepONets for optimal control, we focused on a specific set of assumptions and an open-loop setting. Our investigation was limited to the DeepONet architecture, albeit with some variations in its formulation. However, we see no limitations regarding the use of alternative architectures with likely similar or better performance, as long as they can be trained as physics-informed. This, however, limits the usability of the method, since training a neural operator in a purely physics-informed way is not trivial for complex differential equations. In addition, if the input is not normalized or within a suitable range, physics-informed training becomes difficult, further restricting general applicability. We also restricted the class of admissible controls to smooth functions, which simplifies the analysis and optimization but may not capture discontinuous or bang-bang control strategies. Finally, we assumed well-posed and bounded solution spaces, leaving the treatment of more complex or unbounded systems to future work, as well as more complicated PDEs. Overall, we view our results as a first step toward implementing physics-informed neural operators as such, e.g, without additional architectural modifications, as surrogate models for optimal control. Future studies will be needed to better understand both their limitations and their potential.
ACKNOWLEDGMENT
This work was supported by the Finnish Ministry of Education and Culture’s Pilot for Doctoral Programmes (Pilot project Mathematics of Sensing, Imaging and Modelling).